OpenStudy (anonymous):

A courtyard is shaped like a square with 250-ft-long sides. What is the distance from one corner of the courtyard to the opposite corner? Round to the nearest tenth.

hero (hero):

250sqrt(2)

OpenStudy (mertsj):

You know, sweetheart, you really do need to learn the special right triangles.

sam (.sam.):

350

OpenStudy (anonymous):

353.6

sam (.sam.):

\[\sqrt{250^{2}+250^{2}}\]

OpenStudy (anonymous):

The square can be divided into two triangles with angles 45, 45 and 90 degrees. The equation is therefore: \[\sqrt(250^2 + 250^2) = 353,6\]